[figure: 9.7 cm]

Vortex pinning in electron-doped cuprate superconductor La_{2−x}Ce_{x}CuO_{4}

Y. Zuev^{a}, T.R. Lemberger^{a}, J. Skinta^{a}
T. Greibe^{b}, M. Naito^{b}

a Department of Physics, Ohio State University, Columbus, OH, 43210, U.S.A. b NTT Basic Research Laboratories, Morinosato Wakamiya, Atsugi-shi, Kanagawa 243,

Japan.

We have measured the vortex pinning strength in thin films of electron-doped cuprate La_{2−x}Ce_{x}CuO_{4} for various doping x from x = 0.075 (T_{c} = 25K) to x = 0.15, (T_{c} = 12.5K). Optimal doping for this compound is x ≈ 0.11, at which T_{c} = 29K. Films were prepared by molecular-beam epitaxy method. Structural XRD investigation revealed no doping-induced lattice strain. We use two-coil technique at frequencies from 10 to 100kHz. Superfluid density n_{s}(0) grows rapidly with x on underdoped side and decreases slowly on overdoped side. The Labusch parameter κ, i.e. the spring constant for pinned vortex, shows large variations as doping changes. It increases with x rapidly to a maximum at slight underdoping and then decreases on overdoped sight more rapidly then n_{s}(0). Even at optimal doping κ(0)is two orders of magnitude lower then the highest values observed in YBCO.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : Yuri Zuev

Address for correspondence : Department of Physics, Ohio State University 174 w.18^{th} ave. Columbus, OH,43210 U.S.A.

Email address : zuev.2@osu.edu